OpenStudy (anonymous):

8b+3-10b=-2(b-2)+3 8b+3-10b=-2b-4+3 18b+3=2b-7 -2 -2 16b+3=-7 -3 -3 16b -10 16 16 b= where did I go wrong?

OpenStudy (anonymous):

2nd line/ -4 should be +4

OpenStudy (anonymous):

18b +3=2b+7 -2 -2 16b +3=7 -3-3 16b= 4 16 16 did I do something wrong?

OpenStudy (anonymous):

yeah 8b - 10b should be -2b not 18b

OpenStudy (anonymous):

this problem doesn't have a solution

OpenStudy (anonymous):

okay thanks for your help